Logbook:

Day 59, 08:12:32 hours

Personnel adjustment: As noted in amended entry for Day 58, S arrived yesterday. Full details still unknown. Will debrief ASAP, and further ascertain S's current physical and mental condition.

Communication status: projected date of repair completion on target. SOS transmission expected to resume within planned parameters.

Inventory addition: collared-lemming like mammal (probable closest Genus: Dicrostonyx) garnered through twitch-up snare #2. (Note: Tularemia precautions apply as usual.) TUS #1 rendered inoperable: sapling frozen. Remaining traps serviceable but fruitless.

Inventory depletion: Usual daily fuel rations, small quantity of coffee, sugar. Deemed appropriate to boost morale. S may still be suffering shock or after-effects of subzero temperatures.

Weather/Celestial activity: Storm continues with small breaks; breaks averaging approximately 25 minutes. Sunrise projection for yesterday within four minutes of actual. Chart amended. Sunspot activity minimal at 0700.